OpenStudy (anonymous):

Sketch the given curve, using the following steps. Y(x) = 2√x - x A) Domain B) Intercepts C) Asymptotes D) Intervals of increase and decrease + local min and max. E)Concavity + Inflection point.

OpenStudy (anonymous):

This is what I got: A) Domain: (-∞,0) U (0,∞) B) Y-Int = 0 x-Int = 0 C) Increasing at (1,∞) Decreasing at (-∞,1) Local Min: f(1) = 1 No Local Max f) No concavity + inflection point.

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

is this the y? : \(\large y(x)=2\sqrt{x}-x \)

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

I'm not sure about the local min/max and the concavity bit. Also, when I graphed this in my calculator, it showed a point of intersection at (3.8, 0). So I'm really confused.

OpenStudy (anonymous):

Oh, I forgot to include asymptotes to the list. I got no asymptotes.

OpenStudy (anonymous):

the domain is [0, \(\large \infty \))

OpenStudy (anonymous):

Okay, what about (-∞,0)?

OpenStudy (anonymous):

for the x intercepts, set y=0: \(\large 0=2\sqrt{x}-x\rightarrow x^{\frac{1}{2}}(2-x^{\frac{1}{2}})=0 \) so x=0, x=sqrt2 are your x intercepts.

OpenStudy (anonymous):

y-intercept is at y=0

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

you can't stick in any negative values of x into that function... so the domain is [0, infinity)

OpenStudy (anonymous):

Oh, because of the square root? Okay, got it.

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

no asymptotes....

OpenStudy (anonymous):

Okay, that's what I got too.

OpenStudy (anonymous):

\(\large y(x)=2\sqrt{x}-x \) \(\large y(x)=2x^{\frac{1}{2}}-x \) \(\large y '(x)=2(\frac{1}{2})x^{-\frac{1}{2}}-1\rightarrow 0=2(\frac{1}{2})x^{-\frac{1}{2}}-1 \) \(\large x^{-\frac{1}{2}}-1=0 \) \(\huge \frac{1}{x^{\frac{1}{2}}}=1 \rightarrow x=1\)

OpenStudy (anonymous):

so test the function for increasing/decreasing using the first derivative test on the intervals (0, 1) , (1, infinity)

OpenStudy (anonymous):

Okay, so I got increasing at (1,∞) and decreasing at (0,1)?

OpenStudy (anonymous):

Is that okay?

OpenStudy (anonymous):

you got that backwards.... it should be increasing (0, 1), decreasing (1, infinity)

OpenStudy (anonymous):

Oh, my bad!

OpenStudy (anonymous):

So, there's a local max at at x=1 then?

OpenStudy (anonymous):

zero is part of the domain.. so x=0 is a local min. local max at x=1

OpenStudy (anonymous):

Okay. Also, just wanted to double check, you plug in the values into the original equation to find the local min/max, right?

OpenStudy (anonymous):

you can use the first derivative test.... to the left of x=1, y is increasing, to the right of x=1, y is decreasing... so at x=1 is a local max

OpenStudy (anonymous):

Oh okay. I get it.

OpenStudy (anonymous):

same with x=0... but you can only test on the right side of x=0....

OpenStudy (anonymous):

Gotchya! :)

OpenStudy (anonymous):

So, this is what I got as my second derivative f"(x) = -1/2X^(-3/2). When I equate f"(x) = 0, I get 0, concave downward at x=0, no point of inflection.